A student uses an immersion heater to heat some water in a beaker. The water is heated from 20 °C to 80 °C. The energy supplied to the water is 60.0 kJ. What is the thermal capacity of the water? (Ignore any heat loss
1
Expert's answer
2020-04-13T10:06:23-0400

We know that the equation for heating a substance up looks like this:


"Q=cm\Delta t=C\Delta t."

In our problem, the change in temperature is 60 degrees Celsius, therefore, the thermal capacity will be:

"C=\frac{Q}{\Delta t}=\frac{60\cdot10^3}{60}=1000\text{ J/°C}."